6.042/18.062J Mathematics for Computer Science February 1, 2005 Srini Devadas and Eric Lehman

Recitation Sign-Up

YOUR NAME: YOUR EMAIL:

Each recitation section meets for one hour on both Wednesday and Friday. Attendance is mandatory. Please cross out times for which you have a conflict.

9 AM 10 AM 11 AM noon 1 PM 2 PM 3 PM

Please return this sheet at the end of lecture.